what is the mass, in grams, of 9.08 x 10^24 molecules of methanol (CH3OH)?
1
Expert's answer
2019-09-20T07:36:09-0400

N(CH3OH) = 9.08 x 10^24

m (CH3OH) = n*M = (N/NA)*M = (9.08 x 1024/6.02*1023)*32 g/mol = 482.658 g